A customer has a $100 gift card to a local coffee shop. Suppose the customer, using the card, spends $5 per day at the shop. Which recursive rule represents the card’s remaining balance B, in dollars, after using the card for t days?
Answer by greenestamps(13334)



The instructions are self-contradictory. A recursive rule does not represent the balance after t days; it represents the balance after n+1 days in terms of the balance after n days:

B(n+1) = B(n)-5

That recursive rule says "the balance after n+1 days is $5 less than the balance after n days.

A rule that gives the balance after t days is

B(t) = 100-5t

That is an explicit formula -- not a recursive one.